OpenStudy (anonymous):

Solve the following: You can get 12 color prints for $5.76. How much is this per print?

OpenStudy (amistre64):

since: 12 = 5.76 divide both sides by 12

OpenStudy (anonymous):

juz divide = $0.48

OpenStudy (amistre64):

12/12 = 5.76/12 therefore 1 = 5.76/12 eh?

OpenStudy (anonymous):

right........

OpenStudy (lgbasallote):

arent you missing a variable somewhere there @amistre64

OpenStudy (amistre64):

dont need no variables in an identity ...

OpenStudy (amistre64):

pi/180 needs no variables does it?

OpenStudy (amistre64):

might need "unit measures" to be cogent tho lol

OpenStudy (amistre64):

since: 12 prints = 5.76 dollars etc